OpenStudy (anonymous):

Solve the equation for c and show your work. -2(c + 19) = -52

Parth (parthkohli):

Division of both sides by -2; subtract 19 from both then.

OpenStudy (anonymous):

(c+19)=26 c=26-19 c=7

OpenStudy (anonymous):

Thank You! Can you help me some more please nitz

OpenStudy (anonymous):

ya sure......